40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 23 05:22:51 2011, kiwamu, Update, LSC, PRMI locking : not stable enough 
    Reply  Thu Jun 23 09:46:16 2011, rana, Update, LSC, PRMI locking : not stable enough 
Message ID: 4863     Entry time: Thu Jun 23 05:22:51 2011     Reply to this: 4864
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: PRMI locking : not stable enough 

I was trying to measure the sensing matrix in the PRMI configuration, but basically gave up.

It is mainly because the lock of PRMI wasn't so stable and it didn't stay locked for more than a minute.

It looked like an angular motion fluctuated a lot around 1- 3 Hz. The beam spot on the AS camera moved a lot during the lock.

I have to figure out who is the bad suspension and why.

ELOG V3.1.3-